OpenStudy (anonymous):

A ball of mass 0.120 kg is dropped from rest from a height of 1.25 m. It rebounds from the floor to reach a height of 0.600 m. What impulse was given to the ball by the floor?

OpenStudy (anonymous):

mass (m) = 0.120kg initial velocity (Vo) = 0 came from rest distance traveled (s) = 1.25 m acceleration (g) = 9.8 m/s^2 due to gravity final velocity (Vf) = ? \[Vf ^{2} = 2gs + Vo ^{2}\] \[Vf ^{2} = 2(9.8 m/s ^{2})(1.25 m) + 0\] \[Vf = 7\sqrt{2}\div2 m/s\] \[Vf \approx 4.950 m/s\] \[change \in momentum (\Delta P) = Impulse (J)\] \[\Delta P = m( Vf - Vo )\] \[\Delta P = 0.120kg(7\sqrt{2}\div2 m/s)\] \[\Delta P \approx 0.594 kg m/s\] kg m/s = Ns ( Newton*seconds ) Ns is the common unit of Impulse Impulse is approximately 0.594 Ns, directed upward (opposing the direction of velocity)